Question Hel In a study of pleas and prison sentences tis found that 63% of the subjects studied were sent to prison Among those sent to prison, 40% chose to plead guilty Among those not sent to prison, 35% chose to plead guilty. a. Ifone of the study subjects is randomly selected. ind the probability of getting someone who was not sent to prison b. If a study subject is randomly selected and t is then found that the subject entered a guilty plea, find the probability that this person was not a. I one of the study subjects is randomly selected, find the probability of getting P(Not sent to prison)-A (Type an integer or a decimal.) b. If a study subject is randomly selected and it is then found that the subject entered a guilty plea, find the probability that this person was not sent to prison P(Not sent to prison(Guilty plea) Round to three decimal places as needed) someone who was not sent to prison. Enter your answer in each of the answer boxes Save for Later Carlo S North Carolina State with similar questions
Answer #1

a) P(Not sent to prison) =1-P(Sent to Prison) = 1-0.63 = 0.37

b) P(Guity plea) =P(Sent to prison and plea guilty)+P(not sent to prison and plea guilty) = 0.63*0.40+0.37*0.35 =0.3815

P(Not sent to prison and guilty plea) = 0.37*0.35 = 0.1295

P(Not sent to Prison| Guilty plea) = 0.1295/0.3815 = 0.33945